OpenStudy (anonymous):

Integrate.. (tan(lnx^4))/x

OpenStudy (anonymous):

is it (lnx)^4 or ln(x^4)?

OpenStudy (anonymous):

\[\int\limits \frac{ \tan(lnx ^{4} )}{ x }\]

OpenStudy (anonymous):

scratch that just u substitute for 4 lnx

OpenStudy (anonymous):

^ agreed

OpenStudy (anonymous):

so du =4/x

OpenStudy (anonymous):

\[\int\limits\limits \frac{ \tan(4 \ln x )}{ x }dx\]\[u=4 \ln x\]\[{du \over 4} = {1 \over x}dx\]Replace 1/x dx with du/4

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }\int\limits(\tan(u))\] how do i integrate tan

OpenStudy (anonymous):

\[\int\limits \tan u \space du=-\ln|\cos u| + C\]courtesy of my Diff EQ book

OpenStudy (anonymous):

tan(u)=sin(u)/cos(u)..if you play with that you get what ChmE got

OpenStudy (anonymous):

\[\frac{ -1 }{ 4 }\ln \left| \cos(4lnx) \right|\] got it thanks guys

OpenStudy (anonymous):

yw